OpenStudy (anonymous):

Which is the equation for the line perpendicular to y = -5/3x + 11 1/3 and containing P(-2, 3)? y - 2 = -3/5(x - 3) y = -5/3x + 4 1/3 y = -3/5x + 4 1/5 y = 3/5 + 4 1/5

OpenStudy (anonymous):

@gorv @Battleheroes01

OpenStudy (gorv):

first find the slope of given line on comparing with standard form of line euation y=mx+c m=slope slope of given line= -5/3

OpenStudy (gorv):

let slope of perpendicular line= m so product of slope must = -1 -5/3 * m=-1 m=3/5

OpenStudy (anonymous):

So it's C?